Quiz # 2 for Dr. Z.’s Number Theory Course for Oct. 3, 2013

NAME: (print!) E-MAIL ADDRESS: (print!)

1. (5 points) Use any method to find the proudct-of-prime-powers representation of (3^{3} · 5^{98} · 11^{14} · 43) · (3^{2} · 5^{3} · 11^{12} · 13^{2} · 43^{2}) .

2. (5 points) Use the clever way to find Div(7007).